OpenStudy (anonymous):

determine the quotient and the remainder using the Horner. a. (6x^2+x-8)/(2x-3) b. (3x^3+19x^2+11x+11)/(3x-1)

OpenStudy (emmigrace222):

im doing this with you threw pm :)

OpenStudy (anonymous):

@Ganeshie8, @campbell_st, @primeralph, @ajprincess, @nurali. all of you, can you help me?

OpenStudy (emmigrace222):

really -.-

OpenStudy (emmigrace222):

REALLY -.- ok im sorry but i can help you anymore

OpenStudy (anonymous):

\[ \Large 6 x^2+x-8=(2 x-3) (3 x+5)+7\\ \Large 3 x^3+19 x^2+11 x+11=(3 x-1) \left(x^2+\frac{20 x}{3}+\frac{53}{9}\right)+\frac{152}{9} \] Use the above to check your solution

OpenStudy (emmigrace222):

heyy that person id not online